A tugboat tows a ship at a constant speed of 14.7ft/sec. If the tension in the tow rope remains steady value of 2,000.00lb, calculate the required power output of the tugboats motor in kilowatts.

a.2.27 kw b. 39.89 kw c. 53.45 kw d. 136.05kw



1
Expert's answer
2021-04-20T16:50:26-0400

By definition, power is:


"P = 14.7\dfrac{ft}{s}\cdot 2000lb =29400\dfrac{pound-force}{s}"

One pound-force/s is equal to 1.3558 watts. Thus:


"P = 29400\cdot 1.3558\space W \approx 39861\space W\approx 39.86\space kW"

Answer. b. 39.89 kw